\begin{thebibliography}{00}

\bibitem[\protect\citeauthoryear{Abraham}{1903}]{Abr03}
Abraham, M., 1903,
Prinzipien der Dynamik des Elektrons,
Ann. Phys. (Leipzig), 310, 501
%
\bibitem[\protect\citeauthoryear{Amsler et al.}{2008}]{Amsetal}
Amsler, C., Doser, M., Antonelli, M., \& 170~authors, 2008,
Review of particle physics,
Phys. Lett. B, 667, 1
%
\bibitem[\protect\citeauthoryear{Ashby \& Allan}{1979}]{AshAll}
Ashby, N., \& Allan, D.W., 1979,
Practical implications of relativity for a global coordinate time scale,
Radio Science, 14, 649
%
\bibitem[\protect\citeauthoryear{Ballmer et al.}{2010}]{Baletal}
Ballmer, S., M\'arka, S., \& Shawhan, P., 2010,
Feasibility of measuring the Shapiro time delay over meter-scaled distances,
Class. Quant. Grav., 27, 185018
%
\bibitem[\protect\citeauthoryear{Bauch \& Weyers}{2002}]{BauWey}
Bauch, A., \& Weyers, S., 2002,
New experimental limit on the validity of local position invariance,
\prd, 65, 081101
%Phys. Rev. D,
%
\bibitem[\protect\citeauthoryear{Bersons}{2013}]{Ber13}
Bersons, I., 2013,
Soliton model of the photon,
Lat. J. Phys. Tech. Sci., 50, 60
%
\bibitem[\protect\citeauthoryear{Bersons et al.}{2014}]{Beretal}
Bersons, I., Veilande, R., \& Pirktinsh, A., 2014,
Three-dimen\-sional collinearly propagating solitons,
Physica Scripta, 89, id. 045102
%
\bibitem[\protect\citeauthoryear{Blamont \& Roddier}{1961}]{BlaRod}
Blamont, J.E., \& Roddier, F., 1961,
Precise observation of the profile of the Fraunhofer strontium resonance line.
Evidence for the gravitational red shift on the Sun,
\prl, 7, 437
%
\bibitem[\protect\citeauthoryear{Boonserm et al.}{2005}]{Booetal}
Boonserm, P., Cattoen, C., Faber, T., Visser, M., \& Weinfurtner,~S., 2005,
Effective refractive index tensor for weak-field gravity,
Class. Quant. Grav., 22, 1905
%
\bibitem[\protect\citeauthoryear{Bondi}{1986}]{Bon86}
Bondi, H., 1986,
Relativity theory and gravitation,
Eur. J. Phys., 7, 106
%
\bibitem[\protect\citeauthoryear{Bopp}{1929}]{Bop29}
Bopp, K., (Ed.), 1929,
Fatio de Duillier: De la cause de la pesanteur,
Schriften der Stra{\ss}burger Wiss. Ges. Heidelberg, 10, 19
%
\bibitem[\protect\citeauthoryear{Brault}{1962}]{Bra62}
Brault, J.W., 1962,
The gravitational redshift in the solar spectrum,
PhD Diss., Princeton University
%
\bibitem[\protect\citeauthoryear{Brault}{1963}]{Bra63}
Brault, J.W., 1963,
Gravitational redshift of solar lines,
Bull. Am. Phys. Soc., 8, 28
%
\bibitem[\protect\citeauthoryear{de Broglie}{1923}]{Bro23}
de Broglie, L., 1923,
Ondes et quanta,
Comptes rendus, 177, 507

\bibitem[\protect\citeauthoryear{Cacciani et al.}{2006}]{Cacetal}
Cacciani, A., Briguglio, R., Massa, F., \& Rapex, P., 2006,
Precise measurement of the solar gravitational red shift,
Celest. Mech. Dyn. Astron., 95, 425
%
\bibitem[\protect\citeauthoryear{Chen \& Kantowski}{2008}]{CheKan}
Chen, B., \& Kantowski, R., 2008,
Cosmology with a dark refraction index
\prd, 78, id. 044040
%Phys. Rev. D.
%
\bibitem[\protect\citeauthoryear{Chou et al.}{2010}]{Choetal}
Chou, C.W., Hume, D.B., Rosenband, T., \& Wineland, D.J., 2010,
Optical clocks and relativity,
Science, 329, 1630
%
\bibitem[\protect\citeauthoryear{Cranshaw et al.}{1960}]{Craetal}
Cranshaw, T.E., Schiffer, J.P., \& Whitehead, A.B., 1960,
Measurement of the
gravitational red shift using the M\"ossbauer effect in Fe$^{57}$,
\prl, 4, 163
%
\bibitem[\protect\citeauthoryear{Desloge}{1990}]{Des90}
Desloge, E.A., 1990,
The gravitational red shift in a uniform field,
Am. J. Phys., 58, 856

\bibitem[\protect\citeauthoryear{Dicke}{1960}]{Dic60}
Dicke, R.H., 1960,
E\"otv\"os experiment and the gravitational red shift,
Am. J. Phys., 28, 344
%
\bibitem[\protect\citeauthoryear{Dirac}{1927}]{Dir27}
Dirac, P.A.M., 1927,
The quantum theory of the emission and absorption of radiation,
Proc. Roy. Soc. Lond. Ser. A, 114, 243
%
\bibitem[\protect\citeauthoryear{Dirac}{1951}]{Dir51}
Dirac, P.A.M., 1951,
Is there an {\ae}ther?
\nat, 168, 906
%
\bibitem[\protect\citeauthoryear{Dyson et al.}{1920}]{Dysetal}
Dyson, F.W., Eddington, A.S., \& Davidson, C., 1920,
A determination of the deflection of light by the Sun's
gravitational field, from observations made at the
total eclipse of May 29, 1919,
Phil. Trans. R. astr. Soc. Lond. A, 220, 291
%
\bibitem[\protect\citeauthoryear{Earman \& Glymour}{1980}]{EarGly}
Earman, J., \& Glymour, C., 1980,
The gravitational red shift as a test of general relativity: History
and analysis,
Stud. Hist. Philos. Sci. A, 11, 175
%
\bibitem[\protect\citeauthoryear{Einstein}{1905a}]{Ein05a}
Einstein, A., 1905a,
\"Uber einen die Erzeugung und Verwandlung des Lichtes
betreffenden heuristischen Gesichtspunkt,
Ann. Phys. (Leipzig), 322, 132
%
\bibitem[\protect\citeauthoryear{Einstein}{1905b}]{Ein05b}
Einstein, A., 1905b,
Zur Elektrodynamik bewegter K\"orper,
Ann. Phys. (Leipzig), 322, 891
%
\bibitem[\protect\citeauthoryear{Einstein}{1908}]{Ein08}
Einstein, A., 1908,
\"Uber das Relativit\"atsprinzip und die aus demselben gezogenen Folgerungen,
Jahrbuch der Radioaktivit\"at und Elektronik 1907, 4, 411
%
\bibitem[\protect\citeauthoryear{Einstein}{1911}]{Ein11}
Einstein, A., 1911,
\"Uber den Einflu{\ss} der Schwerkraft auf die Ausbreitung des Lichtes,
Ann. Phys. (Leipzig), 340, 898
%
\bibitem[\protect\citeauthoryear{Einstein}{1912}]{Ein12}
Einstein, A., 1912,
Lichtgeschwindigkeit und Statik des Gravitationsfeldes,
Ann. Phys. (Leipzig), 343, 355
%
\bibitem[\protect\citeauthoryear{Einstein}{1916}]{Ein16}
Einstein, A., 1916,
Die Grundlage der allgemeinen Relati\-vi\-t\"ats\-theorie,
Ann. Phys. (Leipzig), 354, 769
%
\bibitem[\protect\citeauthoryear{Einstein}{1917}]{Ein17}
Einstein, A., 1917,
Zur Quantentheorie der Strahlung,
Phys. Z., XVIII, 121
%
\bibitem[\protect\citeauthoryear{Einstein}{1920}]{Ein20}
Einstein, A., 1920,
\"Ather und Relativit\"ats\-theorie,
Rede gehalten am 5. Mai 1920
an der Reichs-Universit\"at zu Leiden,
Verlag von Julius Springer, Berlin
%
\bibitem[\protect\citeauthoryear{Einstein}{1924}]{Ein24}
Einstein, A., 1924,
\"Uber den \"Ather,
Verhandl. Schweiz. Naturforsch. Gesell., 105, 85
%
\bibitem[\protect\citeauthoryear{Ellis}{2010}]{Ell10}
Ellis, R.S., 2010,
Gravitational lensing: A unique probe of dark matter and dark energy,
Phil. Trans. A: Math. Phys. Eng. Sci., 368, 1914, 967
%
\bibitem[\protect\citeauthoryear{Fatio de Duillier}{1690}]{Fat90}
Fatio de Duillier, N., 1690,
De la cause de la pesanteur,
Not. Rec. Roy. Soc. London, 6, 2 (May 1949), 125
%
\bibitem[\protect\citeauthoryear{Fermi}{1932}]{Fer32}
Fermi, E., 1932,
Quantum theory of radiation,
Rev. Mod. Phys., 4, 87
%
\bibitem[\protect\citeauthoryear{Feynman et al.}{1995}]{Feyetal}
Feynman, R.P., Morinigo, F.B., \& Wagner, W.G., 1995,
(Ed. Brian Hatfield),
Feynman lectures on gravitation,
Reading, MA, Addison-Wesley
%
\bibitem[\protect\citeauthoryear{Flie{\ss}bach}{2006}]{Fli06}
Flie{\ss}bach, T., 2006,
Allgemeine Relativit\"atstheorie,
Elsevier, Spektrum Akademischer Verlag, Heidelberg
%
\bibitem[\protect\citeauthoryear{Gagnebin}{1949}]{Gag49}
Gagnebin, B., 1949,
Introduction: 1. Nicolas Fation de Duillier, 2. La commuication de Fatio de
Duillier, 3. Les manuscrits de Fatio de Duillier
Not. Rec. Roy. Soc. London, 6, 2 (May 1949), 105
%
\bibitem[\protect\citeauthoryear{Goldhaber \& Nieto}{1971}]{GolNie}
Goldhaber, A.S., \& Nieto, M.M., 1971,
Terrestrial and extraterrestrial limits on the photon mass,
Rev. Mod. Phys., 43, 277
%
\bibitem[\protect\citeauthoryear{Granek}{2001}]{Gra01}
Granek, G., 2001,
Einstein's ether: F. Why did Einstein come back to the ether?
Apeiron, 8, 19
%
\bibitem[\protect\citeauthoryear{Hay et al.}{1960}]{Hayetal}
Hay, H.J., Schiffer, J.P., Cranshaw, T.E., \& Egelstaff, P.A., 1960,
Measurement of the red shift in an accelerated system using the M\"ossbauer
effect in Fe$^{57}$,
\prl, 4, 165
%
\bibitem[\protect\citeauthoryear{Hentschel}{1993a}]{Hen93a}
Hentschel, K., 1993a,
The discovery of the redshift of solar Fraunhofer lines
by Rowland and Jewell in Baltimore around 1890,
Hist. Stud. Phys. Biol. Sci. , 23, 2, 219
%
\bibitem[\protect\citeauthoryear{Hentschel}{1993b}]{Hen93b}
Hentschel, K., 1993b,
The  conversion of St.\,John:
A case study on the interplay of theory and experiment,
Sci. Context 6, 1, 137
%
\bibitem[\protect\citeauthoryear{Hentschel}{1996}]{Hen96}
Hentschel, K., 1996,
Measurements of gravitational redshift between 1959 and 1971,
Ann. Sci., 53, 269
%
\bibitem[\protect\citeauthoryear{Jackson}{2006}]{Jac06}
Jackson, J.D., 2006,
Klassische Elektrodynamik, 4.~Aufl.,
Walter de Gruyter, Berlin, New York
%
\bibitem[\protect\citeauthoryear{Jewell}{1896}]{Jew96}
Jewell, L.E., 1896,
The coincidence of solar and metallic lines. A study of the appearence
of lines in the spectra of the electric arc and the Sun,
\apj, 3, 89
%
\bibitem[\protect\citeauthoryear{Kamenov \& Slavov}{1998}]{KamSla}
Kamenov, P., \& Slavov, B., 1998,
The photon as a soliton,
Found. Phys. Lett., 11, 325
%
\bibitem[\protect\citeauthoryear{Kollatschny}{2004}]{Kol04}
Kollatschny, W., 2004,
AGN black hole mass derived from the gravitational redshift in optical lines,
Proc. IAU Symp., 222, 105
%
\bibitem[\protect\citeauthoryear{Kramer et al.}{2006}]{Kraetal}
Kramer, M., Stairs, I.H., Manchester, R.N., \& 12~coauthors, 2006,
Tests of general relativity from timing the double pulsar,
Science, 314, 97
%
\bibitem[\protect\citeauthoryear{Krause \& L\"uders}{1961}]{KraLue}
Krause, I.Y., \& L\"uders, G., 1961,
Experimentelle Pr\"ufung der Relativit\"atstheorie
mit Kern\-re\-sonanz\-absorption,\\
Naturwiss., 48, 34
%
\bibitem[\protect\citeauthoryear{Kutschera \& Zajiczek}{2010}]{KutZaj}
Kutschera, M., \& Zajiczek, W., 2010,
Shapiro effect for relativistic
particles\,--\,Testing general relativity in a new window,
Acta Phys. Polonica~B, 41, 1237
%
\bibitem[\protect\citeauthoryear{L\"ammerzahl}{2009}]{Lae09}
L\"ammerzahl, C., 2009,
What determines the nature of gravity? A phenomenological approach,
\ssr, 148, 501
%Space Sci. Rev.,
%
\bibitem[\protect\citeauthoryear{Landau \& Lifchitz}{1972}]{LanLif}
Landau, L., \& Lifchitz, E., 1972,
Th\'eorie quantique relativiste I,
\'Editions Mir, Moscou
%
\bibitem[\protect\citeauthoryear{von Laue}{1912}]{Lau12}
von Laue, M., 1912,
Zwei Einw\"ande gegen die Relativit\"atstheorie und ihre Widerlegung,
Phys. Z., XIII, 118
%
\bibitem[\protect\citeauthoryear{von Laue}{1920}]{Lau20}
von Laue, M., 1920,
Zur Theorie der Rotverschiebung der Spektrallinien an der Sonne,
Z. Phys., 3, 389
%
\bibitem[\protect\citeauthoryear{von Laue}{1959}]{Lau59}
von Laue, M., 1959,
Geschichte der Physik, 4. erw. Aufl.,
Ullstein Taschenb\"ucher-Verlag, Frankfurt/Main
%
\bibitem[\protect\citeauthoryear{Lewis}{1926}]{Lew26}
Lewis, G.N., 1926,
The conservation of photons,
\nat, 118, 874
%
\bibitem[\protect\citeauthoryear{LoPresto et al.}{1980}]{LoPetal}
LoPresto, J.C., Chapman, R.D., \& Sturgis, E.A., 1980,
Solar gravitational redshift,
\solphys, 66, 245
%Sol. Phys.,
%
\bibitem[\protect\citeauthoryear{LoPresto et al.}{1991}]{LoP91}
LoPresto, J.C., Schrader, C., \& Pierce, A.K., 1991,
Solar gravitational redshift from the infrared oxygen triplet,
\apj, 376, 757
%
\bibitem[\protect\citeauthoryear{Mannheim}{2006}]{Man06}
Mannheim, P.D., 2006,
Alternatives to Dark Matter and Dark Energy,
Prog. Particle Nucl. Phys., 56, 340
%
\bibitem[\protect\citeauthoryear{Meulenberg}{2013}]{Meu13}
Meulenberg, A., 2013,
The photonic soliton,
\procspie, 8832, id. 88320M
%Proc. SPIE,
%
\bibitem[\protect\citeauthoryear{Michelson \& Morley}{1887}]{MicMor}
Michelson, A.A., \& Morley, E.W., 1887,
On the relative motion of the Earth and of the luminiferous ether,
Sidereal Messenger, 6, 306
%
\bibitem[\protect\citeauthoryear{Michelson et al.}{1928}]{Mic28}
Michelson, A.A. Lorentz, H.A., Miller, D.C., Kennedy, R.J., Hedrick, E.R.,
\& Epstein, P.S., 1928,
Conference on the Michelson--Morley experiment held at Mount Wilson,
February, 1927,
\apj, 68, 341
%
\bibitem[\protect\citeauthoryear{Mikhailov}{1959}]{Mik59}
Mikhailov, A.A., 1959,
The deflection of light by the gravitational field of the Sun,
\mnras, 119, 593
%
\bibitem[\protect\citeauthoryear{M{\o}ller}{1957}]{Moe57}
M{\o}ller, C., 1957,
On the possibility of terrestrial tests of the general theory of relativity,
Nuovo Cimento, 6, 381
%
\bibitem[\protect\citeauthoryear{M\"uller et al.}{2010}]{Mueetal}
M\"uller, H., Peters, A., \& Chu, S., 2010,
M\"uller, Peters \& Chu reply,
\nat, 467, E2
%
\bibitem[\protect\citeauthoryear{Negi}{2005}]{Neg05}
Negi, P.S.,2005,
An upper bound on the energy of a gravitationally redshifted
electron-positron annihilation line from the Crab pulsar,
A\&A, 431, 673
%
\bibitem[\protect\citeauthoryear{Nimtz \& Stahlhofen}{2008}]{NimSta}
Nimtz, G., \& Stahlhofen, A.A., 2008,
Universal tunneling time for all fields,
Ann. Phys. (Berlin), 17, 374
%
\bibitem[\protect\citeauthoryear{Ohanian}{1976}]{Oha76}
Ohanian, H.C., 1976,
Gravitation and spacetime,
W.W. Norton, New York
%
\bibitem[\protect\citeauthoryear{Okun}{2000}]{Oku00}
Okun, L.B., 2000,
Photons and static gravity,
Mod. Phys. Lett. A, 15, 1941
%
\bibitem[\protect\citeauthoryear{Okun et al.}{2000}]{Okuetal}
Okun, L.B., Selivanov, K.G., \& Telegdi, V.L., 2000,
On the interpretation of the redshift in a static gravitational field,
Am. J. Phys., 68, 115
%
\bibitem[\protect\citeauthoryear{Pasquini et al.}{2011}]{Pasetal}
Pasquini, L., Melo, C., Chavero, C., Dravins, D., Ludwig, H.-G.,
Bonifacio, P., \& de La Reza, R., 2011,
Gravitational redshifts in main-sequence and giant stars,
A\&A, 526, id.A127
%
\bibitem[\protect\citeauthoryear{Poincar\'e}{1900}]{Poi00}
Poincar\'e, H., 1900,
La th\'eorie de Lorentz et le principe de r\'eaction,
Arch. N\'eerland. Sci. exact. natur., 5, 252
%
\bibitem[\protect\citeauthoryear{Pound}{2000}]{Pou00}
Pound, R.V., 2000,
Weighing photons,
Class. Quantum Grav., 17, 2303
%
\bibitem[\protect\citeauthoryear{Pound \& Rebka}{1959}]{PouReb}
Pound, R.V., \& Rebka, G.A., 1959,
Gravitational red-shift in nuclear resonance,
\prl, 3, 439
%
\bibitem[\protect\citeauthoryear{Pound \& Snider}{1965}]{PouSni}
Pound, R.V., \& Snider, J.L., 1965,
Effect of gravity on gamma radiation,
Phys. Rev., 140, 788
%
\bibitem[\protect\citeauthoryear{Preston}{1875}]{Pre75}
Preston, S.T., 1875,
Physics of the ether,
E. \& F.\,N. Spon, London, New York
%
\bibitem[\protect\citeauthoryear{Quinn et al.}{2014}]{Quietal}
Quinn, T., Speake, C., Parks, H., \& Davis, R., 2014,
The BIPM measurements of the Newtonian constant of gravitation, G.
\prl 112, 039901
%
\bibitem[\protect\citeauthoryear{Randall}{2006}]{Ran06}
Randall, L., 2006,
Verborgene Universen, 4. Aufl.,
S. Fischer Verlag GmbH, Frankfurt/Main
%
\bibitem[\protect\citeauthoryear{Reasenberg et al.}{1979}]{Reaetal}
Reasenberg, R.D., Shapiro, I.I., MacNeil, P E., Goldstein, R.B.,
Breidenthal, J.C., Brenkle, J.P., Cain, D.L., Kaufman, T.M.,
Komarek, T.A., \& Zygielbaum, A.I., 1979,
Viking relativity experiment:
Verification of signal retardation by solar gravity,
\apj, 234, L219
%
\bibitem[\protect\citeauthoryear{Rowland}{1896}]{Row96}
Rowland, H. A., 1896,
On a table of standard wave lengths of the spectral lines,
Memoirs of the American Academy of Arts and Sciences, New Series, 12, 2, 101
%
\bibitem[\protect\citeauthoryear{Schiff}{1960}]{Sch60}
Schiff, L.I., 1960,
On experimental tests of the general theory of relativity,
Am. J. Phys., 28, 340
%
\bibitem[\protect\citeauthoryear{Schr\"oder}{1990}]{Sch90}
Schr\"oder, W., 1990,
Ein Beitrag zur fr\"uhen Diskussion um den \"Ather
und die Einsteinsche Relativit\"atstheorie,
Ann. Phys., 7. Folge, 326, Heft 6, 475
%
\bibitem[\protect\citeauthoryear{Scott}{2015}]{Sco15}
Scott, R.B., 2015,
Teaching the gravitational redshift:
Lessons from the history and philosophy of physics,
J. Phys.: Conf. Ser., 600,1, id. 012055
%
\bibitem[\protect\citeauthoryear{Shapiro}{1964}]{Sha64}
Shapiro, I.I., 1964,
Fourth test of general relativity,
\prl, 13, 789
%
\bibitem[\protect\citeauthoryear{Shapiro et al.}{1971}]{Sha71}
Shapiro, I.I.,  Ash, M.E., Ingalls, R.P., Smith, W.B.,
Campbell,~D.B., Dyce, R.B., Jurgens, R.F., \& Pettengill, G.H., 1971,
Fourth test of general relativity: New radar result,
\prl, 26, 1132
%
\bibitem[\protect\citeauthoryear{Sinha \& Samuel}{2011}]{SinSam}
Sinha, S., \& Samuel, J., 2011,
Atom interferometry and the gravitational redshift,
Class. Quantum Grav., 28, 145018-1
%
\bibitem[\protect\citeauthoryear{Snider}{1970}]{Sni70}
Snider, J.L., 1970,
Atomic-beam study of the solar 7699~\AA \\ potassium line and the solar
gravitational red-shift,
\solphys, 12 352
%
\bibitem[\protect\citeauthoryear{Snider}{1972}]{Sni72}
Snider, J.L., 1972,
New measurement of the solar gravitational red shift,
\prl, 28, 853
%
\bibitem[\protect\citeauthoryear{Soldner}{1804}]{Sol04}
Soldner, J., 1804,
Ueber die Ablenkung eines Lichtstrals von seiner geradlinigen
Bewegung durch die Attraktion eines Weltk\"orpers, an welchem er nahe vorbei
geht,
Berliner Astron. Jahrb. 1804, 161
%
\bibitem[\protect\citeauthoryear{Sommerfeld}{1978}]{Som78}
Sommerfeld, A., 1978,
Optik,
Verlag Harri Deutsch, Thun, Frankfurt/Main
%
\bibitem[\protect\citeauthoryear{Sotiriou et al.}{2008}]{Sotetal}
Sotiriou, T.P., Liberati, S., \& Faraoni, V., 2008,
Theory of gravitation theories: A no-progress report,
Int. J. Mod. Phys. D, 17, 399
%
\bibitem[\protect\citeauthoryear{St.\,John}{1928}]{StJ28}
St.\,John, C.E., 1928,
Evidence for the gravitational displacement of lines in the solar spectrum
predicted by Einstein's theory,
\apj, 67, 195
%
\bibitem[\protect\citeauthoryear{Straumann}{2004}]{Str04}
Straumann, N., 2004,
General relativity with applications to astrophysics,
Springer-Verlag, Berlin, Heidelberg, New York
%
\bibitem[\protect\citeauthoryear{Takeda \& Ueno}{2012}]{TakUen}
Takeda, Y., \& Ueno, S., 2012,
Detection of gravitational redshift on the solar disk by using
iodine-cell technique,
\solphys, 281, 551
%
\bibitem[\protect\citeauthoryear{Turyshev}{2013}]{Tur13}
Turyshev, S.G., 2013,
Testing fundamental gravitation in space,
Nucl. Phys. B (Proc. Suppl.) 243, 197
%
\bibitem[\protect\citeauthoryear{Vessot et al.}{1980}]{Vesetal}
Vessot, R.F.C., Levine, M.W., Mattison, E.M., Blomberg, E.L., Hoffman, T.E.,
Nystrom, G.U., Farrel, B.F., Decher, R., Eby,~P.B., \& Baugher, C.R., 1980,
Test of relativistic gravitation with a space-borne hydrogen maser,
\prl, 45, 2081
%
\bibitem[\protect\citeauthoryear{Vigier}{1991}]{Vig91}
Vigier, J.-P., 1991,
Explicit mathematical construction of relativistic nonlinear de Broglie waves
described by three-dimensional (wave and electromagnetic) solitons
``piloted'' (controlled) by corresponding solutions of associated linear
Klein-Gordon and Schr\"odinger equations,
Found. Phys., 21, 125
%
\bibitem[\protect\citeauthoryear{Wiechert}{1911}]{Wie11}
Wiechert, E., 1911,
Relativit\"atsprinzip und \"Ather,
Phys. Z., 12, 689 \& 737
%
\bibitem[\protect\citeauthoryear{Wilhelm \& Dwivedi}{2013}]{WilDwi13}
Wilhelm, K., \& Dwivedi, B.N., 2013,
Gravity, massive particles, photons and Shapiro delay,
(Paper~2),
\apss, 343, 145
%
\bibitem[\protect\citeauthoryear{Wilhelm \& Dwivedi}{2014}]{WilDwi14}
Wilhelm, K., \& Dwivedi, B.N., 2014,
On the gravitational redshift,
\na, 31, 8
%
\bibitem[\protect\citeauthoryear{Wilhelm et al.}{2013a}]{WiDwWi}
Wilhelm, K., Dwivedi, B.N., \& Wilhelm, H., 2013a,
An impact model of the electrostatic force: Coulomb's law re-visited,
arXiv1403.1489
%
\bibitem[\protect\citeauthoryear{Wilhelm et al.}{2013b}]{WiWiDw}
Wilhelm, K., Wilhelm, H., \& Dwivedi, B.N., 2013b,
An impact model of Newton's law of gravitation,
(Paper~1),
\apss, 343, 135
%
\bibitem[\protect\citeauthoryear{Will}{1974}]{Wil74}
Will, C.M., 1974,
Gravitational red-shift measurements as tests of nonmetric theories of gravity,
Phys. Rev. D, 10, 2330
%
\bibitem[\protect\citeauthoryear{Will}{2006}]{Wil06}
Will, C.M., 2006,
The confrontation between general relativity and experiment,
Living Rev. Relativity 9, 3
%
\bibitem[\protect\citeauthoryear{Wolf et al.}{2010}]{Woletal}
Wolf, P., Blanchet, L., Bord\'e, C.J., Reynaud, S., Salomon, C., \&
Cohen-Tannoudji, C., 2010,
tom gravimeters and gravitational redshift,
\nat, 467, E1
%
\bibitem[\protect\citeauthoryear{Ye \& Lin}{2008}]{YeLin}
Ye, X.-H., \& Lin, Q., 2008,
Gravitational lensing analysed by the graded refractive index of a vacuum,
J. Opt. A: Pure Appl. Opt., 10, 7, id. 075001
%
\end{thebibliography}

\documentclass{aastex}
\usepackage{spr-astr-addons}
\usepackage{url}

\RequirePackage{color}
\renewcommand{\vec}[1]{\mbox{\boldmath $#1$}}

\newcommand{\emaila}{wilhelm@mps.mpg.de}
\newcommand{\emailb}{bnd.app@iitbhu.ac.in}
\renewcommand{\vec}[1]{\mbox{\boldmath $#1$}}
\def\arcsec{\hbox{$^{\prime\prime}$}}

\newcommand*\Del{\mathrm{\Delta}}                 
\newcommand{\rmd}{ {\ \mathrm d} }
\newcommand{\uvec}[1]{\hat{\vec #1}}
\newcommand{\m}{ {\ \mathrm m} }
\newcommand{\s}{ {\ \mathrm s} }
\newcommand{\kg}{ {\ \mathrm {kg}} }

\newcommand{\grad}{{\bf \nabla } }
\begin{document}

\title{Gravitational redshift and the vacuum index of refraction}
\shorttitle{Gravitational redshift}
\shortauthors{K. Wilhelm, B.N. Dwivedi}

\author{Klaus Wilhelm}
\affil{Max-Planck-Institut f\"ur Son\-nen\-sy\-stem\-for\-schung
(MPS),\\ Justus-von-Liebig-Weg 3, 37077 G\"ottingen, Germany\\ \emaila}

\and

\author{Bhola N. Dwivedi}
\affil{Department of Physics, Indian Institute of Technology
(Banaras Hindu University), Varanasi-221005, India\\ \emailb}

Last updated on \today

\vspace{1cm}

\begin{abstract}
A physical process of the gravitational redshift was described in an earlier
paper \citep{WilDwi14} that did not require any information for the emitting
atom neither on the local gravitational potential~$U$ nor on the speed of
light~$c$. Although it could be shown that the correct energy shift of the
emitted photon resulted from energy and momentum conservation principles and
the speed of light at the emission site, it was not obvious how this speed is
controlled by the gravitational potential. The aim of this paper is to
describe a physical process that can accomplish this control. We determine the
local speed of light~$c$ by deducing a gravitational index of
refraction~$n_{\rm G}$ as a function of the potential~$U$ assuming a specific
aether model, in which photons propagate as solitons. Even though an atom
cannot locally sense the gravitational potential~$U$
\citep[cf.][]{Mueetal}, the gravitational redshift will nevertheless be
determined by~$U$ \citep[cf.][]{Woletal}\,---\,mediated by the local speed of
light~$c$.
\end{abstract}

\keywords{Gravitation, impact model, aether, gravitational index of
refraction}

\section{Introduction} 
\label{s:introd}
\label{sec:intro}

The study of the gravitational redshift, predicted for solar radiation by
\citet{Ein08}, is still an important subject in modern physics and
astrophysics \citep[e.g.,][]{Kol04,Neg05,Lae09,Choetal,Pasetal,Tur13}.
The displacement of metallic lines to the violet observed in the laboratory
in comparison with the corresponding solar lines had first been noted by
\citet{Row96} and \citet{Jew96} \citep[cf.][]{Hen93a}.
Measurements of the small gravitational redshift of solar
spectral lines are inherently difficult, because many processes in the
atmosphere of the Sun can influence the spectrum. In particular, the high
speeds of the emitting plasmas lead to line shifts due to the classical Doppler
effect \citep[cf.][]{Hen93b}.
Nevertheless, early observations confirmed Einstein's prediction in general
\citep{StJ28,BlaRod,Bra62,Bra63,Sni70} \citep[cf.][]{Hen96}.
Improved observational techniques \citep[e.g.,][]{LoPetal,Cacetal,TakUen},
have established a shift of solar lines of
%
\begin{eqnarray}
c_0\,\frac{\Del \lambda}{\lambda} \approx 600\m\s^{-1} ~,
\label{shift_600}
\end{eqnarray}
where $c_0 = 299\,792\,458\m\s^{-1}$ is the speed of light in vacuum
remote from any masses and $\lambda$ the wavelength of the electromagnetic
radiation.

The gravitational potential~$U$ at a distance~$r$ from a spherical body with
mass~$M$ is constraint in the weak-field approximation for non-relativistic
cases \citep[cf.][]{LanLif} by
%
\begin{eqnarray}
- 1  \ll \frac{U}{c^2_0} = - \frac{G_{\rm N}\,M}{c^2_0~r} \le 0 ~,
\label{potential}
\end{eqnarray}
where $G_{\rm N} = 6.67554(16) \times 10^{-11}\kg^{-1}\m^3\s^{-2}$
is Newton's constant of gravity \citep{Quietal}.
A definition of a reference potential in line with
Eq.~(\ref{potential}) is $U_0 = 0$ for $r = \infty$.

In an attempt to describe the physical process(es) that lead to the
gravitational redshift,
\citet{Woletal} and \citet{Mueetal} disagreed on whether the frequency
of an atomic clock is sensitive to the
gravitational potential~$U$ (according to Wolf et al.) or, as suggested by
M\"uller et al., to the local gravity field $\vec{g} = \grad U$.
Support for the first alternative can be found in many publications
\citep[e.g.,][]{Ein08,Lau20,Sch60,Wil74,Okuetal,SinSam}, but it is, indeed,
not obvious how an atom can locally sense the gravitational potential~$U$.

Experiments on Earth~\citep{PouReb,Craetal,KraLue,PouSni},
in space~\citep{Vesetal,BauWey} and in the Sun-Earth
system~\citep{StJ28,BlaRod,Bra62,Bra63,Sni72,LoP91,Cacetal,TakUen} have,
however, quantitatively confirmed in the static weak field approximation a
relative frequency
shift of
%
\begin{eqnarray}
\frac{\nu - \nu_0}{\nu_0} = \frac{\Del \nu}{\nu_0}
\approx \frac{\Del U}{c^2_0} = \frac{U - U_0}{c^2_0}~,
\label{shift}
\end{eqnarray}
where $\nu_0 = c_0/\lambda_0$ is the frequency of the radiation emitted by a
certain transition at~$U_0$ and $\nu$ the observed frequency there, if the
emission caused by the same transition had occurred at a potential~$U$.

In addition to the redshift, the deflection of light near gravitational
centres is of fundamental importance. For a close solar fly-by
\citet{Sol04} and
\citet{Ein11} obtained $0.87\arcsec$ under the assumption
that radiation would be affected in the same way as matter.
\emph{Twice} this value was then derived in the framework of the General
Theory of Relativity \citep[GTR,][]{Ein16}\footnote{It is of interest in
the context of this paper that Einstein employed Huygens' Principle in his
calculation of the deflection.},
and later by \citet{Sch60} using the equivalence principle and STR.
The high value was confirmed
during the total solar eclipse in 1919 for the first time \citep[]{Dysetal}.
This and later observations have been summarized by \citet{Mik59} and
combined to a mean value of
approximately 2\arcsec.

\section{Graviton interactions} 
\label{sec:graviton}

A model of gravitational interactions based on a modified impact concept
has been proposed for massive bodies \citep[][Paper~1]{WiWiDw},
and the difficulties of the old theory proposed by
Nicolas \citet{Fat90} \citep[cf.][]{Bop29,Gag49} have been
considered in the light of the Special Theory of Relativity
\citep[STR,][]{Ein05a} and the non-local behaviour of virtual
particles \citep[cf.][]{NimSta}. The basic idea is that impacting
\emph{gravitons}\,---\,originally named \emph{quadrupoles}\,---\,with no mass
and a speed $c_0$ are absorbed by massive particles and re-emitted with
reduced energy~$T^-_{\rm G}$ according to
%
\begin{eqnarray}
T^-_{\rm G} = T_{\rm G}\,(1 - Y) ~,
\label{Eq:reduced_energy}
\end{eqnarray}
where $T_{\rm G}$ is the energy of a graviton in the background flux and
$0 < Y \ll 1$.
A spherically symmetric emission of a liberated graviton with a reduction
parameter~$Y$ had been assumed in Paper~1. Further studies
have, however, indicated that an anti-parallel emission with respect to the
incoming graviton is more appropriate, because conflicts with the energy and
momentum conversation principles in closed systems can be avoided by the
second choice.
Newton's law of gravitation could be explained with this model,
however, a secular mass increase of matter was a consequence
of its application. This poses the question of how the interaction of gravity
with photons can be understood, since the photon mass is in all likelihood
zero.\footnote{A zero mass of photons follows from the STR and a speed of light
in vacuum~$c_0$ constant for all frequencies. \citet{Ein05b} used
,,Lichtquant'' for a quantum of electromagnetic radiation;
the term ``photon'' was introduced by \citet{Lew26}.
With various methods the photon mass could be constrained to
$m_\nu < 10^{-49}~\kg$ \citep{GolNie,Amsetal}.}
If the mass of a photon is indeed zero, the interaction process must be
different.
An initial attempt at solving that problem has been made in Paper~2
\citep{WilDwi13} and is summarized here under the assumption of an
anti-parallel re-emission, both for massive particles and photons.

A physical process will then be outlined that provides information on
the gravitational potential~$U$ at the site of a photon emission. This aspect
had not been covered in our earlier paper on the gravitational redshift
\citep{WilDwi14}.

Interactions between massive bodies have been treated in Paper~1 with an
absorption rate of \emph{half} the intrinsic de Broglie frequency~$m\,c^2_0/h$
for a mass~$m$ \citep[cf.][]{Bro23}, because \emph{two} virtual gravitons have
to be emitted for one interaction, whereas in Paper~2 it is assumed that a
photon causes a reflection with an interaction rate of $\nu = E_\nu/h$ with
Planck's constant~$h$. The momentum transfer to a photon will thus be twice
as high as to a massive body with a mass equivalent to $E_\nu/c^2_0$.

If we apply the momentum conservation principle to photon-graviton pairs in
the same way as to photons \citep[cf.][]{LanLif}, we can write after a
reflection of $\vec {p}_{\rm G}$
%
\begin{eqnarray}
\vec {p}_\nu + 2\,\vec {p}_{\rm G} =
\vec {p}^*_\nu
\label{Eq:momentum}
\end{eqnarray}
with $|{\vec p}_{\rm G}| = p_{\rm G} = T_{\rm G}/c_0$.

We assume, applying Eq.~(\ref{Eq:momentum}) with
$p_{\rm G} \ll p_\nu = |\vec {p}_\nu|$, that
under the influence of a gravitational centre relevant
interactions occur on opposite sides of a photon with $p_{\rm G}$ and
$p_{\rm G}\,(1 - Y)$ transferring a net momentum of $2\,Y\,p_{\rm G}$. Note,
in this context, that the Doppler effect can only operate for interactions of
photons with massive bodies \citep[cf.][]{Fer32,Som78}.
Consequently, there will be no energy change of the photon, because both
gravitons are reflected with constant energies under these
conditions, and we can write for a pair of interactions
%
\begin{eqnarray}
E_\nu = |\vec p_\nu|\,c = |\vec p_\nu + 2\,Y\vec p_{\rm G}|\,c' =
|\vec p'_\nu|\,c' = E'_\nu ~,
\label{Eq:photon_energy}
\end{eqnarray}
where $\vec p'_\nu$ is the photon momentum after the events. If $\vec p_\nu$
and a component of $2\,Y\vec p_{\rm G}$ are pointing in the same direction,
it is $c' < c$, the speed is reduced; an antiparallel direction leads to
$c' > c$. Note that this could, however, not result in $c' > c_0$, because
$c = c_0$ can only be attained in a region with
an isotropic distribution of gravitons with a momentum of $p_{\rm G}$,
i.e. with a gravitational potential~$U_0 = 0$.

The momentum $\vec p_\nu$ of a photon radially approaching a gravitational
centre will be treated in line with Eq.~(6) in
Sect.~2 of Paper~2 for massive bodies, however, with twice the interaction
rate (valid for photons as explained above).
Since we know from observations that the deflection
of light during a close fly-by at the Sun is very small\,--\,to simplify the
calculations, we only treat this configuration\,--\,the
momentum variation caused by the weak and static
gravitational interaction is also very small.
The momentum change rate of the photon can then be approximated by
%
\begin{eqnarray}
\frac{\Del {\vec p}_\nu}{\Del t_M} \approx
2\,G_{\rm N}\,M\,\frac{\uvec r}{r^2}\,\frac{p_\nu}{c_0} ~,
\label{Eq:transfer_rate}
\end{eqnarray}
where $M$ the mass of the gravitational centre, $r = |\vec r|$ the
distance of the photon from the centre, and the position vector of
the photon is $r\,\uvec{r}$ with a  unit vector $\uvec{r}$.
The small deflection angle also allows us to approximate
the actual path by a straight line and use $x \approx c_0\,t_M$
along an $x$~axis.
The normalized momentum variation along the trajectory then is
%
\begin{eqnarray}
-\frac{c_0}{p_\nu}\left(\frac{\Del \vec{p}_\nu}{\Del t_M}\right)_x =
\frac{c_0}{p_\nu}\,\frac{\Del p_\nu}{\Del t_M}\cos \vartheta
\approx 2\,G_{\rm N}\,M\,\frac{x}{r^3}~.
\label{Eq:x_component}
\end{eqnarray}
The corresponding component perpendicular to the trajectory is
%
\begin{eqnarray}
-\frac{c_0}{p_\nu}\left(\frac{\Del \vec{p}_\nu}{\Del t_M}\right)_y =
\frac{c_0}{p_\nu}\,\frac{\Del p_\nu}{\Del t_M}\,\sin \vartheta
\approx 2\,G_{\rm N}\,M\frac{R}{r^3}~,
\label{Eq:y_component}
\end{eqnarray}
where $R$ is the impact parameter of the trajectory.
Integration of Eq.~(\ref{Eq:x_component}) over $t_M$ from $-\infty$ to $x/c_0$
yields
%
\begin{eqnarray}
\frac{1}{p_\nu}\,[\rmd \vec{p}_\nu(r)]_x \approx
\frac{2\,G_{\rm N}\,M}{c_0^2\,r} =
\frac{2\,G_{\rm N}\,M}{c_0^2\,\sqrt{R^2 + x^2}}~.
\label{Eq:x_integrated}
\end{eqnarray}

If we apply Eq.~(\ref{Eq:photon_energy}) to a photon approaching the
mass~$M$ along the $x$~axis
starting from infinity with $E_\nu = p_\nu\,c_0$, and
considering that the $y$~component in Eq.~(\ref{Eq:x_component}) is much smaller
than the x component in Eq.~(\ref{Eq:y_component}) for $x \gg R$,
the photon speed~$c(r)$ as a function of $r$
can be determined from
%
\begin{eqnarray}
p_\nu\,c_0 \approx \{p_\nu + [\rmd \vec{p}_\nu(r)]_x\}\,c(r)~.
\label{Eq:reduced_c}
\end{eqnarray}
Division by $p_\nu\,c_0$ then gives with Eq.~(\ref{Eq:x_integrated})
%
\begin{eqnarray}
\frac{1}{[n_{\rm G}(r)]_x} = \frac{c(r)}{c_0} \approx
1 - \frac{2\,G_{\rm N}\,M}{c_0^2~r} = 1 + \frac{2\,U(r)}{c_0^2}
\label{Eq:refraction}
\end{eqnarray}
as a good approximation of the inverse gravitational index of refraction
along the $x$~axis. The same index has been obtained albeit with different
arguments, e.g., by \citet{Booetal,YeLin}. The resulting speed of light is in
agreement with evaluations by \citet{Sch60}, for a radial
propagation\footnote{\citet{Ein12} states
explicitly that the speed at a certain location is not dependent on the
direction of the propagation.} in a central gravitational field, and
\citet{Oku00}\,---\,calculated on the basis of the standard Schwarzschild
metric. A decrease of the speed of light near the Sun, consistent with
Eq.~(\ref{Eq:refraction}), is not only supported by the predicted and
subsequently observed Shapiro delay
\citep{Sha64,Reaetal,Sha71,Kraetal,Baletal,KutZaj},
but also indirectly by the deflection of light \citep{Dysetal}.

\section{Gravitational redshift}
\label{sec:grav_red}

Since Einstein discussed the gravitational redshift and published conflicting
statements regarding this effect, the confusion could still not be cleared up
consistently \citep[cf., e.g.,][]{Man06,Sotetal}. In most of his publications
Einstein defined clocks as atomic clocks. Initially he assumed that the
oscillation of an atom corresponding to a spectral line might be an
intra-atomic process, the frequency of which would be determined by the atom
alone \citep{Ein08,Ein11}. \citet{Sco15} also felt that the equivalence
principle and the notion of an ideal clock running independently of
acceleration suggest that such clocks are unaffected by gravity. \citet{Ein16}
later concluded that clocks would slow down near gravitational centres thus
causing a redshift.

The question whether the gravitational redshift is caused by the emission
process (Case~a) or during the transmission phase (Case~b) is nevertheless
still a matter of recent debates. Proponents of (a) are, e.g.,
\citet{Moe57,Des90,Sch60,Craetal,Oha76,EarGly,Oku00,Okuetal} and of
(b): \citet{Hayetal,Feyetal,Str04,Fli06,Ran06,Wil06}.

There is general agreement on the observational and experimental facts and
most of the arguments are formally consistent with them, but different
physical processes or mathematical concepts are considered. In particular, it
is surprising that the same team of experimenters, albeit with different first
authors (Cranshaw et al. and Hay et al.) published different views on the
process of the Pound--Rebka--Experiment. \citet{PouSni} and \citet{Pou00}
pointed out, however,
that this experiment could not distinguish between the two options, because
the invariance of the velocity of the radiation had not been demonstrated.
\citet{Bon86} and \citet{Dic60} also left the question open. In many cases,
the confusion results from the unclear definitions of clocks and times as
detailed, for instance, by \citet{AshAll} and \citet{Oku00}.

\citet{Ein17} emphasized that for an elementary emission process not only
the energy exchange, but also the momentum transfer is of importance
\citep[cf., as well][]{Poi00,Abr03,Fer32}. Taking these considerations into
account, \citet{WilDwi14} formulated a photon emission process at
a gravitational potential~$U$ assuming that:
\begin{enumerate}
\item [(1)] The atom cannot sense the potential~$U$, in line with the
original proposal by \citet{Ein08,Ein11}, and initially emits the same
energy~$\Del E_0$ at $U > 0$ and $U_0 = 0$.
\item [(2)] It also cannot directly sense the speed of light at the location with a
potential~$U$. The initial momentum thus is $p_0 = \Del E_0/c_0$.
\item [(3)] As the local speed of light is, however, $c(U) \ne c_0$, a photon
having an energy of $\Del E_0$ and a momentum $p_0$ is not able to propagate.
The necessary adjustments of the photon energy and momentum as well as the
corresponding atomic quantities then lead in the \emph{interaction region} to
a redshift consistent with $h \nu = \Del E_0\,(1 + U/c^2_0)$ and observations.
\end{enumerate}

As outlined in Sect.~\ref{sec:graviton}, there is general
agreement in the literature that the local speed of light is
%
\begin{eqnarray}
c(U) \approx c_0 \left(1 + \frac{2\,U}{c^2_0}\right)
\label{Eq:local_speed}
\end{eqnarray}
in line with Eq.~(\ref{Eq:refraction}). It has, however, to be noted that
in Sect.~\ref{sec:graviton} the speed~$c(U)$ was obtained for a photon
propagating from $U_0$ to $U$, and, therefore, the physical process which
controls the speed of newly emitted photons is not established. An attempt
to do that will be made in the next section.

\section{An aether model with photons as solitons}
\label{sec:aether}
Before we suggest a specific aether model, a few statements on the aether
concept in general should be mentioned. Following \citet{MicMor} famous
experiment, \citet{Ein05a,Ein08} concluded that the concept of a light aether
as carrier of the electric and magnetic forces is not consistent with the STR.
In response to critical remarks by \citet{Wie11}, cf. \citet{Sch90} for
Wiechert's support of the aether, \citet{Lau12} wrote that the existence
of an aether is not a physical, but a philosophical problem, but later
differentiated between the physical world and its mathematical formulation.
A four-dimensional `world' is only a valuable mathematical trick; deeper
insight, which some people want to see behind it, is not involved \citep{Lau59}.

In contrast to his earlier statements, Einstein said at the end of a speech
in Leiden that according to the GTR a space without aether cannot be
conceived \citep{Ein20}; and even more detailed: Thus one could instead of
talking about `aether' as well discuss the `physical properties of space'.
In theoretical physics we cannot do without aether, i.e., a continuum endowed
with physical properties \citep{Ein24}.
\citet{Mic28} confessed at a meeting in Pa\-sa\-de\-na in the
presence of H.A. Lorentz that he clings a liitle to the aether; and
\citet{Dir51} wrote in a letter to Nature that there are good reasons for
postulating an {\ae}ther.

\citet{WiDwWi} proposed an impact model for the electrostatic force
based on massless \emph{dipoles}.
The vacuum is thought to be permeated by these dipoles that are, in the absence of
electromagnetic or gravitational disturbances, oriented and directed randomly
propagating along their dipole axis with a speed of~$c_0$. There is little or
no interaction among them. Note that such electric dipoles have no mean
interaction energy, even in the classical theory \citep[see, e.g.,][]{Jac06}.
We suggest to identify the dipole distribution with an aether. This is very
similar to the conclusion of \citet{Pre75}:
\begin{enumerate} \item[ ]
``[...] first, that the normal
state of the component particles of the ether is a state of motion; second,
that this motion of the particles takes place in straight lines; and third,
that this motion takes place towards every possible direction.''
\end{enumerate}
Einstein's aether mentioned above may, however, be more related to the
gravitational interactions \citep[cf.][]{Gra01}. In this case, we have to
consider the graviton distribution as another component of the aether.

If we assume that an individual dipole interacts with gravitons in the same
way as photons, see Eq.~(\ref{Eq:photon_energy}), according to
%
\begin{eqnarray}
E_{\rm D} = |\vec p_{\rm D}|\,c = |\vec p_{\rm D} + 2\,Y\vec p_{\rm G}|\,c' =
|\vec p'_{\rm D}|\,c' = E'_{\rm D} ~,
\label{Eq:dipole_energy}
\end{eqnarray}
where $E_{\rm D}$ and $\vec p_{\rm D}$ refer to the energy and momentum of a
dipole. We can then modify Eqs.~(\ref{Eq:transfer_rate}) to (\ref{Eq:reduced_c})
by changing $\nu$ to D and find that Eqs.~(\ref{Eq:refraction}) and
(\ref{Eq:local_speed} ) are also valid for dipoles with a speed of $c_0$ for
$U_0 = 0$. One exception from Preston's ``ether'' is that dipoles can, according
to a modified Eq.~(\ref{Eq:y_component}), be deflected by graviton
interactions.

Considering that many suggestions have been made to describe photons as
solitons \citep[e.g.,][]{Dir27,Vig91,KamSla,Meu13,Ber13,Beretal}, we
also propose that a photon is a soliton propagating in the dipole
aether with a speed of~$c(U)$, cf., Eq.~(\ref{Eq:local_speed}), controlled by
the dipoles moving in the direction of propagation of the photon.
The dipole distribution thus determines the gravitational index of
refraction, cf. Eq.~(\ref{Eq:refraction}), and consequently the speed of
light~$c(U)$ at the potential~$U$. This solves the problem formulated
at the end of Sect.~\ref{sec:grav_red} and might be relevant for other
phenomena, such as gravitational lensing and the cosmological redshift
\citep[cf., e.g.,][]{Ell10,CheKan}.

We will further assume that the dipoles constituting a photon will have turned
the orientation of their axes to a direction perpendicular to the photon
velocity vector. This avoids any electrostatic interactions
during emission and absorption processes of photons, and will probably also be
required by their polarization effects.

\section{Discussion and Conclusion}
\label{sec:concl}
Our aim was to identify a physical process that leads to a speed~$c(U)$ of
photons controlled by the gravitational potential~$U$. This could be achieved
by postulating an aether model with moving dipoles, in which a
gravitational index of refraction $n_{\rm G}(U) = c_0/c(U)$ regulates the
emission and propagation of photons as required by energy and momentum
conservation principles. The emission process thus follows Steps~(1) to (3)
in Sect.~\ref{sec:grav_red}, where the local speed of light is given by the
gravitational index of refraction~$n$. In this sense, the statement that an
atom cannot detect the potential~$U$ by \citet{Mueetal} is correct; the local
gravity field~$\vec{g}$, however, is not controlling the emission process.

A photon will be emitted by an atom with appropriate energy and momentum
values, because the local speed of light requires an adjustment of the
momentum. This occurs in the interaction region between the atom and its
environment as outlined in Step~(3) of Sect.~\ref{sec:grav_red}. A receiver
of the same type next to the emitter would also not be able to determine the
potential either, because the energy and momentum restrictions apply for the
absorption process as well.


\acknowledgments
This research has made extensive use of the
Smithsonian Astrophysical Observatory (SAO)/National Aeronautics and Space
Administration (NASA) Astrophysics Data System (ADS).
Administrative support has been provided by the Max-Planck-Institute for Solar
System Research and the Indian Institute of Technology (Banaras Hindu
University).

\end{document}